Factoring Trinomials by Trial and Error
b) y=-2x2 + 28x - 26 with the factorial form
c) y=-9x2 + 72x + 81 with the factorial form

Found 2 solutions by Nate, CrazyMan Jr.:
Answer by Nate(3500)

You can easily do it by trail and error, but you can also look at my example if you get fed up with it.
2x^2 - 14x + 24
2(x^2 - 7x + 12)
2(x^2 - 4x - 3x + 12)
2((x^2 - 4x) + (-3x + 12))
2(x(x - 4) - 3(x - 4))
2(x - 3)(x - 4)
(2x - 6)(x - 4)

Answer by CrazyMan Jr.(21)

b)-2x^2+28x-26 ~~~ m*n = -13 ~~~ 13*-1
-2(^2+14x-13) ~~~ m-n = 14 ~~~ 13+1
c)-9x^2+72x+81 ~~~ m*n = 9 ~~~ 9*1
-9(x^2+8x+9) ~~~ m-n = 8 ~~~ 9-1

J-Man